Calculate how many grams of aluminum are contained in 510 grams Al O
1
Expert's answer
2020-03-24T14:02:29-0400

The aluminum (II) oxide is an interesting substance detected in stellar absorption spectra. Since the molar mass of aluminum is 27 g/mol and the molar mass of atomic oxygen is 16 g/mol, we can calculate what portion of mass accounts for aluminum:


"\omega(\text{Al})=\frac{\Mu(\text{Al})}{\Mu(\text{AlO})}."

Therefore, the mass of aluminum contained in 510 g of the oxide is


"m(\text{Al})=\omega(\text{Al})m(\text{AlO})=\frac{\Mu(\text{Al})}{\Mu(\text{AlO})}m(\text{AlO})=\\
\space\\
=\frac{27}{27+16}\cdot510=320\text{ g}."